# Near-Optimal Adaptive Transforms for LLM Quantization

### v1

@@ A.2 Laplacian Distribution
Let $X_{1},\dots,X_{d}$ be zero-mean Laplacian random variables with diversity $b_{1},\dots,b_{d}$. Denote $b_{\star}=\max_{k}b_{k}=\sqrt{\frac{1}{2}\max_{k}\mathbb{E}X_{k}^{2}}$. For $t\geq 0$, $\mathbb{P}\left(\left|X_{k}\right|\geq t\right)=\exp\left(-b_{k}^{-1}t\right)\leq\exp\left(-b_{\star}^{-1}t\right)$. Using Lemma 5,
$\ \mathbb{E}\max_{k}X_{k}^{2}=2\int_{0}^{\infty}t\ \mathbb{P}\left(\max_{k}\left|X_{k}\right|\geq t\right)\mathrm{d}t=2\int_{0}^{\infty}t\ \mathbb{P}\left(\bigcup_{k=1}^{d}\left\{\left|X_{k}\right|\geq t\right\}\right)\mathrm{d}t$ (30)
@@
$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+2d\int_{t_{\star}}^{\infty}t\exp\left(-b_{\star}^{-1}t\right)\mathrm{d}t=\left[t^{2}\right]_{0}^{t_{\star}}+2d\left[-b_{\star}\left(t+b_{\star}\right)\exp\left(-b_{\star}^{-1}t\right)\right]_{t_{\star}}^{\infty}$
@@
Splitting at $t_{\star}=b\ln d$ gives

### v2

@@ A.5.2 Laplacian Distribution
Let $X_{1},\dots,X_{d}$ be zero-mean Laplacian random variables with diversity $b_{1},\dots,b_{d}$. Denote
$b_{\star}=\max_{k}b_{k}=\max_{k}\sqrt{\frac{1}{2}\mathbb{E}X_{k}^{2}}=\sqrt{\frac{1}{2}\max_{k}\mathbb{E}X_{k}^{2}}.$ (93)
For $t\geq 0$,
$\mathbb{P}\left(\left|X_{k}\right|\geq t\right)=\exp\left(-b_{k}^{-1}t\right)\leq\exp\left(-b_{\star}^{-1}t\right).$ (94)
Using Lemma 4,
$\mathbb{E}\max_{k}X_{k}^{2}=$ $2\int_{0}^{\infty}t\ \mathbb{P}\left(\max_{k}\left|X_{k}\right|\geq t\right)\mathrm{d}t$ (95)
$=$ $2\int_{0}^{\infty}t\ \mathbb{P}\left(\bigcup_{k=1}^{d}\left\{\left|X_{k}\right|\geq t\right\}\right)\mathrm{d}t$
@@
$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+2d\int_{t_{\star}}^{\infty}t\exp\left(-b_{\star}^{-1}t\right)\mathrm{d}t$
$=$ $\ \left[t^{2}\right]_{0}^{t_{\star}}+2d\left[-b_{\star}\left(t+b_{\star}\right)\exp\left(-b_{\star}^{-1}t\right)\right]_{t_{\star}}^{\infty}$
@@
Splitting at $t_{\star}=b_{\star}\ln d$ gives
@@
Appendix B Further Experimental Results
In this section, we report additional evaluations conducted across the Llama 3 and Qwen 3 family models in Tables˜4, 8, 3, 9, 6 and 7. For the Qwen models, the thinking mode is disabled across all experiments. The evaluation results are obtained using emulated (“fake”) quantization, where quantized values are represented in the BF16/FP16 format. Results are aggregated over multiple random seeds.
Table 3: KL-divergence for Qwen3-8B measured on a slice of C4 dataset, with calibration performed on FineWeb dataset. Results (lower is better) are averaged over multiple random seeds. H denotes the Hadamard transform.
Model Format Quantization Transform KL
Qwen3-8B NVFP4 GPTQ H 0.069518
RTN WUSH 0.065747
GPTQ WUSH 0.054646
MXFP4 GPTQ H 0.093119
RTN WUSH 0.093739
GPTQ WUSH 0.077909
Table 4: W4A4 accuracy results on the LM Evaluation Harness for different models with RTN, GPTQ quantizations, and identity (I), Hadamard (H), WUSH transforms. For the Llama model, the MMLU-CoT metric is reported in the MMLU column.
Model Format Quantization Transform MMLU GSM8K HellaSwag WinoGrande Average Recovery
Llama-3.2-3B-Instruct BF16 - - 64.43 78.01 73.42 70.09 71.49 100.0
NVFP4 RTN I 61.07 72.01 70.90 66.77 67.69 94.68
H 59.91 64.82 69.77 65.59 65.02 90.95
WUSH 59.91 71.24 70.96 67.25 67.34 94.19
GPTQ H 60.22 70.91 71.06 66.57 67.19 93.99
WUSH 60.17 72.71 71.26 67.32 67.87 94.93
MXFP4 RTN I 56.81 60.80 67.30 64.56 62.37 87.24
H 55.99 61.11 68.36 64.09 62.39 87.27
WUSH 59.54 69.49 70.20 66.47 66.43 92.92
GPTQ H 60.24 68.84 69.58 66.08 66.18 92.58
WUSH 59.39 68.26 69.88 67.07 66.15 92.53
Qwen3-8B BF16 - - 72.98 90.90 75.52 70.56 77.49 100.0
NVFP4 RTN I 70.47 88.40 74.44 69.53 75.71 97.70
H 70.19 86.35 73.02 68.11 74.42 96.04
WUSH 70.86 89.92 74.84 70.02 76.42 98.61
GPTQ H 70.78 89.11 74.34 69.25 75.87 97.91
WUSH 71.03 89.56 74.88 70.32 76.45 98.66
MXFP4 RTN I 67.69 84.23 71.24 67.40 72.64 93.74
H 67.53 87.04 71.48 67.64 73.42 94.75
WUSH 69.96 86.90 73.87 68.92 74.91 96.68
GPTQ H 69.58 88.28 73.75 69.25 75.22 97.07
WUSH 70.40 88.80 74.36 69.19 75.69 97.68
Qwen3-14B BF16 - - 77.18 91.96 79.84 74.27 80.81 100.0
NVFP4 RTN I 75.20 90.22 77.38 73.32 79.03 97.80
H 74.98 92.04 77.76 72.38 79.29 98.12
WUSH 75.43 91.78 78.82 73.18 79.80 98.76
MXFP4 RTN I 72.92 90.22 76.68 71.51 77.83 96.31
H 73.13 87.41 76.29 71.67 77.13 95.44
WUSH 74.56 91.87 78.18 73.15 79.44 98.30
Qwen3-32B BF16 - - 80.74 92.12 83.96 76.95 83.44 100.0
NVFP4 RTN I 78.92 90.14 82.60 76.64 82.08 98.37
H 79.48 91.81 82.51 74.98 82.20 98.51
WUSH 79.56 92.72 83.14 75.79 82.80 99.24
MXFP4 RTN I 76.82 75.28 81.32 74.74 77.04 92.33
H 78.63 92.57 81.87 75.53 82.15 98.45
WUSH 78.70 90.83 82.63 75.87 82.01 98.28
Figure 6: Comparison of different transforms on Qwen3-14B for both NVFP4 and MXFP4 RTN quantization on Platinum Benchmarks. The top table shows accuracy results across the individual benchmark tasks, while the bottom plot shows the average accuracy scores together with their standard deviations for each transform.
Figure 7: Comparison of different transforms on Qwen3-32B for both NVFP4 and MXFP4 RTN quantization on Platinum Benchmarks. The top table shows accuracy results across the individual benchmark tasks, while the bottom plot shows the average accuracy scores together with their standard deviations for each transform.
Figure 8: Comparison of different transforms on Llama-3.2-3B-Instruct for both NVFP4 and MXFP4 RTN quantization on Platinum Benchmarks. The top table shows accuracy results across the individual benchmark tasks, while the bottom plot shows the average accuracy scores together with their standard deviations for each transform.
Figure 9: Comparison of different transforms on Llama-3.1-8B-Instruct for both NVFP4 and MXFP4 RTN quantization on Platinum Benchmarks. The top table shows accuracy results across the individual benchmark tasks, while the bottom plot shows the average accuracy scores together with their standard deviations for each transform.